Farmer Fred wants to fence in a rectangular pasture that is bordered on one side by a river. He plans to use 7200 feet of fence to enclose the other three sides. What should the length of the pasture be if the enclosed area is to be a maximum?
Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

maximum area would be a square with 4 equal sides.
we need three sides for the fencing
7200/3= 2400 per side

Answer by Alan3354(69443)

Farmer Fred wants to fence in a rectangular pasture that is bordered on one side by a river. He plans to use 7200 feet of fence to enclose the other three sides. What should the length of the pasture be if the enclosed area is to be a maximum?
--------------
If he fences all 4 sides, the max is a square, area = 1800^2 = 3240000 sq ft.
----
Fencing 3 sides, make the width 1/2 the length --> 1800*3600 = 2x the area of the square.
--------------------
I can work it out. If you need it email via the Thank You note.